OpenStudy (gabbyalicorn):

Which answer shows this equation in standard form? 7 – 3(x – y) = –5x + 2 A. –8x + 3y = –5 B. –2x – 3y = 5 C. 2x + 3y = –5 D. 2x – 3y = –5 so, im guessing after we do 7 – 3(x – y) = –5x + 2 its will turn out to be something involved with distributive property.

OpenStudy (gabbyalicorn):

What i got so far

OpenStudy (campbell_st):

well standard from is normally Ax + By = C or Ax + By + C = 0 so you need to distribute and collect like terms

OpenStudy (campbell_st):

well that's not good Id start with 7 - 3x + 3y = -5x + 2 I did this by distributing the 3 now you can collect like terms...

OpenStudy (gabbyalicorn):

I'm confused on what operations to do for them...I gave it my best shot. @campbell_st

OpenStudy (gabbyalicorn):

:3 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you can't go from `7 – 3(x – y) = –5x + 2` to `4(x – y) = –5x + 2` you have to distribute first, then combine like terms

OpenStudy (gabbyalicorn):

I thought she already distributed when campbell said " I did this by distributing the 3"

OpenStudy (fibonaccichick666):

May I interject something that may help here?

jimthompson5910 (jim_thompson5910):

go ahead

OpenStudy (fibonaccichick666):

So do you know the order of operations gabby?

OpenStudy (gabbyalicorn):

Yes. PEMDAS

OpenStudy (fibonaccichick666):

Good, so, when you are trying to get a variable by itself, you actually can use the reverse order of ops as a guide as to what steps to take. I call it SadMep (look familiar?)

OpenStudy (fibonaccichick666):

So, in this situation, you want to have only x and ys on the left. yea?

OpenStudy (gabbyalicorn):

Yes...

OpenStudy (fibonaccichick666):

ok, so this one has parentheses right?

OpenStudy (gabbyalicorn):

yes.

OpenStudy (fibonaccichick666):

ah, sorry, I want to follow sadmep. not do multiples, ignore the parentheses thing

OpenStudy (fibonaccichick666):

ok, so do we have any x's or y's on the right?

OpenStudy (gabbyalicorn):

Yes. Only 5x.

OpenStudy (fibonaccichick666):

alright, and is it being S or A?

OpenStudy (fibonaccichick666):

We are following SadMep

OpenStudy (gabbyalicorn):

A

OpenStudy (fibonaccichick666):

good, so how do we undo addition?

OpenStudy (gabbyalicorn):

With subtraction.

OpenStudy (fibonaccichick666):

yes, subtraction from Both sides though yea?

OpenStudy (gabbyalicorn):

Yes.

OpenStudy (gabbyalicorn):

So,

OpenStudy (fibonaccichick666):

so if you do that subtraction, we get \[7 – 3(x – y) = –5x + 2\\7 – 3(x – y) -(-5x)= –5x -(-5x)+ 2\\7-3(x-y)-(-5x)=2\] Do you agree?

OpenStudy (gabbyalicorn):

Why do we still have 5x. I thought it got cancelled out.

OpenStudy (fibonaccichick666):

ah, not quite. So 1. you are subtracting -5x from both sides. 2. both sides means both the left side of the equals sign and the right side of the equals sign

OpenStudy (fibonaccichick666):

we just want to move the \(\color\red{-}5x\) to the other side. that was our goal.

OpenStudy (fibonaccichick666):

very important to keep that minus sign too

OpenStudy (gabbyalicorn):

Hmm, okay. By doing that we subtract....

OpenStudy (fibonaccichick666):

well, subtracting a negative is the same as adding the positive. It cancels out

OpenStudy (gabbyalicorn):

Oh. :3

OpenStudy (fibonaccichick666):

so when you do these problems. Write them out as I did above. Use full 'sentences'

OpenStudy (fibonaccichick666):

it will help minimize confusion

OpenStudy (fibonaccichick666):

so we are here: \[7−3(x−y)−(−5x)=2\]

OpenStudy (fibonaccichick666):

now using SadMep to decide what you should do first, can you tell me the next step?

OpenStudy (gabbyalicorn):

Well since we already do subtraction, we add now...but there is nothing to add o-o

OpenStudy (fibonaccichick666):

well, not quite, you actually restart the algorithm every time. So remember our goal of only having x and y terms on the left? Do we have any items that are not x or y over there currently?

OpenStudy (gabbyalicorn):

Yes. 7 & 2.

OpenStudy (fibonaccichick666):

the two is on the left side?

OpenStudy (gabbyalicorn):

No, right...my bad.

OpenStudy (fibonaccichick666):

it's cool, so that 7 though what is it doing? we don't want it there, we need it over on the right hand side. How can we get it there?

OpenStudy (gabbyalicorn):

By doing 5x+2-7?

OpenStudy (fibonaccichick666):

uhm, that doesn't make sense. Remember we want to use full sentences. so you are here:\[7−3(x−y)−(−5x)=2\]

OpenStudy (fibonaccichick666):

what do you need to to to both sides in order to move the 7 to the right side?

OpenStudy (fibonaccichick666):

why did you tuck that little -7 by the 3?

OpenStudy (gabbyalicorn):

I wanted to squeeze the -7 in but not make it overlap the parenthesis or the x

OpenStudy (fibonaccichick666):

So the way you are currently doing it, you changed the problem to: \[-3-7(x-y)-(-5x)=2\] That is actually a completely different problem remember you need to do it to both sides. That means once from the left hand side of the equals sign and once from the right.

OpenStudy (gabbyalicorn):

Well... this would be my only other guess

OpenStudy (fibonaccichick666):

and that would be correct

OpenStudy (gabbyalicorn):

:D Hazaa!

OpenStudy (fibonaccichick666):

now, simplify what you can on each respective side using PEMDAS.

OpenStudy (fibonaccichick666):

tell me what you get

OpenStudy (gabbyalicorn):

I was thinking that I could use like terms and do x- 5x but there's probably no way.

OpenStudy (fibonaccichick666):

so you dropped the minus from the minus minus 5x again. Right hand side looks good. Now, this is where wayyyyyyy up at the top I stepped in

OpenStudy (fibonaccichick666):

This is the distributive property part. To which, I let @campbell_st or @jim_thompson5910 complete their thoughts.

OpenStudy (gabbyalicorn):

Okay, thanks for the help so far. :)

OpenStudy (fibonaccichick666):

so you are here guys: \[-3(x-y)-(-5x)=-5\]

OpenStudy (fibonaccichick666):

np :)

Directrix (directrix):

Standard Form: the standard form of a line is in the form Ax + By = C where A is a positive integer, and B, and C are integers. http://jwilson.coe.uga.edu/emt668/emat6680.2002/jackson/chapter%205%20lesson%20plan/day6.html

OpenStudy (gabbyalicorn):

I was told that earlier by FibonacciChick666, but I don't get how that will help me solve the problem.

OpenStudy (fibonaccichick666):

(it is the template for how you want to make the problem look) Your goal is to have 3 terms total an x term with a coefficient, a y term with a coefficient, and a term that is just a number.

OpenStudy (gabbyalicorn):

okay.

jimthompson5910 (jim_thompson5910):

if the notation `Ax + By = C` confuses you, then think of standard form as this (in the drawing below)